5-METHOXY-N-ACETYL TRYPTOPHAN METHYL AMIDE: THE DRAMATIC EFFECTS OF A REMOTE SUBSTITUTION AND A SINGLE WATER MOLECULE ON ITS CONFORMATIONAL PREFERENCES

\underline{JASPER R. CLARKSON}, BRIAN C. DIAN, and TIMOTHY S. ZWIER, Department of Chemistry, Purdue University, West Lafayette, IN 47907.

Following earlier studies in our group on the conformational preferences of N- acetyl tryptophan methyl amide (NATMA) and N-acetyl tryptophan amide (NATA), we have modified NATMA by substitution of a methoxy group for hydrogen in the 5- position on the indole ring. We present results from a variety of techniques, including LIF, fluorescence-dip infrared spectroscopy (FDIRS), UV-UV hole- burning spectroscopy, and resonant two-photon ionization (R2PI) which characterize the infrared and ultraviolet spectra of this molecule and its water-containing complexes. While NATMA and NATA have three and two significantly populated conformations, respectively, the LIF spectrum of 5- methoxy-NATMA can be accounted for in its entirety by a single conformation. The infrared spectrum of this conformer is consistent with a C5 conformation, in which the dipeptide backbone is in an extended conformation, with no intramolecular H-bond present. Ab initio and density functional theory calculations are used to assign the structure observed, and to understand how the methoxy group restricts the conformations of the dipeptide backbone. When a single water molecule is attached to 5-methoxy-NATMA, a single conformation is observed once again, with water bound to the carboxy amide carbonyl group. However, in this case, the 5-methoxy-NATMA ’solute’ is a C7_{eq} conformation, in which the dipeptide backbone has an intramolecular H-bond. The reasons for this change in conformational preference will be discussed, as will preliminary studies of larger 5-methoxy NATMA-(water)_{n} clusters.